Builtrite had sales of $700,000 and COGS of $290,000. In addition, operating expenses were calculated at 25% of sales. Interest expense was based on $100,000 of bonds outstanding with an interest rate of 7%. Builtrite also received dividends of $40,000 and paid out common stock dividends of $25,000 to its stockholders. A long-term capital gain of $55,000 was realized during the year along with a capital loss of $45,000
Based on the above information, answer the following 4 questions:

What is Builtrite’s taxable income?

Based on their taxable income, what is Builtrite’s tax liability?

Builtrite has $7,000 in interest expense, how much does this interest expense cost Builtrite after taxes?

If Builtrite had experienced a long-term capital loss of $60,000 (instead of the $45,000 long-term capital loss stated in the problem), and still had the $55,000 long-term capital gain stated in the problem, which of the following is correct:

(This problem is not related to the above problem)

Last year Builtrite had retained earnings of $160,000. This year, Builtrite had true net profits after taxesof $65,000 which includes common stock dividends received of $20,000. Builtite also paid a preferred dividend of $15,000. What is Builtrite’s new level of retained earnings?   
Answer #1

1)The taxable income is calculated as shown in the below table:

Sales 700000 Given
Cost of Goods Sold -290000 Given
Gross Profit 410000 Sales - Cost of Goods sold
Operating expenses -175000 25% of sales
EBIT 235000 Gross profit - Operating expenses
Interest expense -7000 7% of 100,000
Dividends received 40000 Given
Net Long term Gain 10000 Capital gain of 55000 - Capital loss of 45000
Taxable Income 278,000

2)For the following calculation tax rate is flat 21% as per the federal tax law.

Builtrite’s taxable income=278000*21%=$58,380

3) Interest expense cost after taxes=7000*(100-21)%=$5,530

4)Net long-term capital loss=60000-55000=(5000)

Earlier there was a net long-term capital gain of $10,000(55000-45000),this is not the case now because long-term capital loss increases to $60,000 resulting in $0(entire $55000 capital gain set-offed against 60000 capital loss) taxable capital gain.

So taxable income would decrease by $10000

AS PER HOMEWORKLIB RULES I'VE ANSWERED THE FIRST 4 PARTS OF THE QUESTION.KINDLY POST THE SAME QUESTION ONCE AGAIN ASKING TO SOLVE THE REMAINING PORTION. PLEASE DON'T GIVE ME A NEGATIVE RATING FOR THIS REASON.IN CASE OF ANY DOUBTS FEEL FREE TO COMMENT BELOW

THANK YOU